# Dynamics of the sine-Gordon equation on tadpole graphs

### v1

@@ 5.1 Morse index for $\mathcal{L}_{Z}$
        Suppose that $g_{{1}}\equiv 0$: then $f_{{1}}(-L)=f_{{1}}(L)=0$, $f^{\prime}_{{1}}(-L)=f^{\prime}_{{1}}(L)$, and so $f_{1}$ will be odd (see step 1 in item A) above). Now, our single-lobe profile $\phi_{1}$ satisfies $\mathcal{L}_{1}\phi^{\prime}_{1}=0$, $\phi^{\prime}$ is odd, $\phi^{\prime}(x)>0$ for $x\in[-L,0)$, thus since $\lambda_{1}<0$ we obtain from the Sturm oscillation theorem that there is $r\in(-L,0)$ such that $\phi^{\prime}(r)=0$, it which is a contradiction. Then, $g_{1}$ is non-trivial.
@@
        Now suppose $g_{1}(0+)=0$: we consider the odd-extension $g_{1,\mathrm{odd}}\in H^{2}(\mathbb{R})$ of $g_{1}$ on all the line and the extension operator $\widetilde{\mathcal{L}}$ in (5.17) on the $\delta$-interaction domains $D_{\delta,\gamma}$ in (5.18). Then, $g_{1,\mathrm{odd}}\in D_{\delta,\gamma}$ for any $\gamma$ and by Perron-Frobenius property (Theorem A.8) we will obtain $n(\widetilde{\mathcal{L}})\geqq 2$, it which is a contradiction because $n(\widetilde{\mathcal{L}})\leqq 1$ for all $\gamma$ (see item (A)(2) above).
@@
    Suppose $\lambda_{0}$ is a double eigenvalue: $\lambda_{0}=\lambda_{1}$. Then by Perron-Frobenius property in item (I)(A) there holds $\xi_{\lambda_{0}},\zeta_{\lambda_{0}}>0$. Now, for $(f_{1},g_{1})\in D_{Z,0}$ an associated eigenfunction for $\lambda_{1}$, we have $(f_{1},g_{1})\perp(\xi_{\lambda_{0}},\zeta_{\lambda_{0}})$. Next, by item (I)(B), Lemma 5.1 and $\lambda_{0}=\eta_{0}$, being $\eta_{0}$ the first eigenvalue (simple) for (PBP) in (5.6), we have that $f_{1}=r\xi_{\lambda_{0}}$ and $g_{1}=s\zeta_{\lambda_{0}}$ with $r=s\neq 0$ by continuity at zero. Hence, we arrive to a contradiction from the orthogonality property of the eigenfunctions. This finishes the proof of the Theorem.
@@
[…]
@@
    Notice that by Theorem 5.8 and item (A) in the proof of Theorem 5.6, we obtain that the unique negative eigenvalue for $(\mathcal{L}_{Z},D_{Z})$ coincides with the smallest negative eigenvalue for the eigenvalues problems (PBP) and ($\delta$BP) in (5.6).
@@
In what follows we exactly determine which single-lobe kink state $(\phi_{1,k},\phi_{2,a})\in D_{Z}$, characterized by Propositions 4.7 and 4.9, will induce that the Morse index of $\mathcal{L}_{Z}$ is equal to one. As we will see the value of $L$, $k$, $c_{i}$ and the sign of $Z$ play a fundamental role.
@@
    with $a=a(k)$ determined by formula in (4.11). Then for $H(k)$ defined in (4.17) we have:
@@
    with $a=a(k)$ determined by formula in (4.11). Then, for $H(k)$ defined in (4.17) we have the following:
@@ 5.2 Trivial kernel for $\mathcal{L}_{Z}$
Let $Z\in(-\infty,\frac{2}{\pi c_{2}})$ be an admissible strength value to obtain a single-lobe kink state $(\phi_{1,k},\psi_{2,a})\in D_{Z}$, given by Propositions 4.7 and 4.9. Then, under the restrictions on $L,c_{1},c_{2},k$ in Theorem 5.10 and 5.12, we obtain $n(\mathcal{L}_{Z})=1$ and $\ker(\mathcal{L}_{Z})=\{\bf{0}\}$. In the case of a degenerate single-lobe kink state $(\pi,\phi_{2,0})\in D_{Z}$ with $Z=\frac{2}{\pi c_{2}}$, we also have that $\ker(\mathcal{L}_{Z})$ is trivial and $n(\mathcal{L}_{Z})=1$.
@@
    Suppose that $c=0$. This implies that $h\equiv 0$ and hence $f$ satisfies $\mathcal{L}_{1}f=0$ with Dirichlet-periodic boundary conditions on $[-L,L]$, $f(L)=f(-L)=0$ and $f^{\prime}(L)=f^{\prime}(-L)$. Suppose that $f\neq 0$:
@@
      Suppose that $f$ is even: then from oscillation results of the Floquet theory, we need to have that zero it is not the first eigenvalue for $\mathcal{L}_{1}$. Therefore, $f$ needs to change of sign and the number of zeros of $f$ on $[-L,L)$ is even. Hence $f$ has an odd number of zeros on $(-L,L)$ (as $f(-L)=0$). Then, $f(0)=0$ and $f^{\prime}(0)=0$. Hence, $f\equiv 0$.
@@
    Suppose that $\alpha_{a}=-Z$ with $Z\leqq 0$. Then $c=0$ and so by the previous item (I) it follows that $\ker(\mathcal{L})=\{\bf{0}\}$. Suppose that $c\neq 0$; without loss of generality we can assume that $c<0$. Hence, $h(x)>0$ on $[0,\infty)$ and by the splitting eigenvalue result in Lemma 5.1 we obtain that $(f,h)$ satisfies the separated eigenvalues problems (PBP)-($\delta$BP) in (5.6) with $\lambda=0$. The $\delta$-interaction condition (also by hypotheses) implies
@@
    Next, we denote by $\{\eta_{n}\}$, $n\in\mathbb{N}_{0}$, the set of eigenvalues for the periodic-problem associated to $\mathcal{L}_{1}$ on $[-L,L]$, and by $\{\mu_{n}\}$, $n\in\mathbb{N}_{0}$, the set of eigenvalues for the Dirichlet-problem associated to $\mathcal{L}_{1}$ on $[-L,L]$. Then, by classical Sturm-Liouville theory (see Theorem 4.8.1 in [52]) we have that $\eta_{0}$ is simple and, in particular, we have the following distribution of eigenvalues,
@@
      Suppose that $0>\mu_{1}$. We know that $\mathcal{L}_{1}\phi_{1}^{\prime}=0$, $\phi^{\prime}_{1}$ is odd and strictly decreasing on $[-L,0]$. Now, since the eigenfunction $\chi$ associated to $\mu_{1}$ is odd and $\chi(-L)=\chi(0)=0$, it follows from the Sturm Comparison Theorem (cf. [51, 52]) that $\phi^{\prime}_{1}$ needs to have one zero on $(-L,0)$. But this is impossible and we have a contradiction.
@@
      From the previous items (a) and (b) it follows that $\mu_{1}>0$ and, by (5.24) and Lemma 5.1 ($\mathcal{L}_{1}f=0$), we need to have $\eta_{1}=0$ and that the zero is simple. Moreover, the eigenfunction $f$ is even (for $f$ being odd, we have $f(L)=-f(-L)=-f(L)$ and so $h(0)=f(L)=0$), which is a contradiction.
    Next, by Floquet theory, $f$ has exactly two different zeros $-\beta,\beta$ ($\beta>0$) on $(-L,L)$. Therefore, $f(0)<0$, $f(\pm L)=h(0)>0$). Now, we consider the Wronskian function (constant) of $f$ and $\phi^{\prime}_{1}$ (we recall $\mathcal{L}_{1}f=\mathcal{L}_{1}\phi^{\prime}_{1}=0$),
@@
The statements $(i)-(iii)$ are immediate from the explicit formula for $\alpha_{a}$. This finishes the proof. ∎
@@
Under the restrictions on $Z,L,c_{1},c_{2}$ and $k$ given by Propositions 4.7 and 4.9 and by Theorems 5.10, 5.12 and 5.14, every member of the smooth family of single-lobe kink states, $k\to(\phi_{1,k},\phi_{2,a(k)})\in D_{Z}$, is spectrally and nonlinearly unstable for the sine-Gordon model (1.4) on a tadpole graph. The degenerate single-lobe kink state $(\pi,\phi_{2,0})\in D_{Z}$, with $Z=\frac{2}{\pi c_{2}}$ and
@@ Acknowledgements
J. Angulo was supported in part by Universal/CNPq project and by CNPq/Brazil Grant. The work of R. G. Plaza was partially supported by DGAPA-UNAM, program PAPIIT, grant IN-104922. R. G. Plaza is grateful to the IME-USP for their hospitality and to the FAPESP, Brazil (processo 2024/15816-0) for their financial support during a research stay in November 2024 when this work was carried out.
@@ Appendix A Appendix
In this section we formulate some tools from the extension theory of symmetric operators by Krein and von Neumann which are suitable for our needs (see, for instance, Naimark [37, 38] and Reed and Simon [44] for further information). In particular, we establish for (~$(\widetilde{\mathcal{L}},D_{\delta,\gamma})_{\gamma\in\mathbb{R}}$ in (5.17) and (5.18), a Perron-Frobenius result to be used in the accurate Morse index estimate $n(\mathcal{L}_{Z})\leqq 1$ for $(\mathcal{L}_{Z},D_{Z})$ in (5.1) and (1.7). We show, for completeness in the exposition, the general estimate $n(\mathcal{L}_{Z})\leqq 2$ using extension theory.
@@ A.1 Extension theory for the Laplacian operator on a tadpole graph
The following three results from the extension theory of symmetric operators are classical and can be found in [44].
@@
By completeness in the exposition, we establish a extension theory for the Laplacian operator
$-\Delta=\Big{(}-\frac{d^{2}}{dx^{2}}\Big{)}_{\mathbb{e}\in\mathbb{E}}$ (A.3)
on a tadpole graph $\mathcal{G}$ with structure represented by the set $\mathbb{E}=(-L,L)\cup(L,+\infty)$ and a single vertex $\nu=L$, such that it induces the one-parameter family of self-adjoint operators $(-\Delta,D_{Z})_{Z\in\mathbb{R}}$ in (5.1) and (1.7). With these notations we have the following result.
Theorem A.4.
The Schrödinger type operator $-\Delta$ in (A.3) on $L^{2}(\mathcal{G})$, with domain
$D(-\Delta)$ =$=\{(\phi,\psi)\in H^{2}(\mathcal{G}):\phi(-L)=\phi(L)=\psi(L)=0,\;\;\phi^{\prime}(L)-\phi^{\prime}(-L)=\psi^{\prime}(L)\}$ (A.4)
is a densely defined symmetric operator with deficiency indices $n_{\pm}(-\Delta)=1$. Therefore, $(-\Delta,D(-\Delta))$ has a one-parameter family of self-adjoint extensions defined by $(-\Delta,D_{Z})_{Z\in\mathbb{R}}$ with
$D_{Z}=\{(f,g)\in H^{2}(\mathcal{G}):f(L)=f(-L)=g(L),\;\text{and},\;f^{\prime}(L)-f^{\prime}(-L)=g^{\prime}(L)+Zg(L)\}.$ (A.5)
Proof.
The symmetric property of $(-\Delta,D(-\Delta))$ is immediate. Since, $C_{c}^{\infty}(-L,L)\bigoplus C_{c}^{\infty}(L,+\infty)\subset D(-\Delta)$ we obtain the density property of $D(-\Delta)$ in $L^{2}(\mathcal{G})$. Now, it is not difficult to see that the adjoint operator $(-\Delta^{*},D(-\Delta^{*}))$ of $(-\Delta,D(-\Delta))$ is given by
$-\Delta^{*}=-\Delta,\quad D(-\Delta^{*})=\{(u,v)\in H^{2}(\mathcal{G}):u(-L)=u(L)=v(L)\}.$ (A.6)
Next, the deficiency subspaces $\mathcal{D}_{\pm}=\ker(-\Delta^{*}\mp i)$ have dimension one. Indeed, by using the definition of $D(-\Delta^{*})$ in (A.6) we can see that $\mathcal{D}_{\pm}=\mathrm{span}\{(f_{\pm i},g_{\pm i})\}$ with
[…]
and,
[…]
Thus, by Proposition A.3, the formulae in (A.7)-(A.8), and the von-Neumann decomposition in Theorem A.1, we can deduce that $(-\Delta,D(-\Delta))$ has a one-parameter family of self-adjoint extensions $(-\Delta,D_{Z})_{Z\in\mathbb{R}}$ with $D_{Z}$ defined in (A.5). This finishes the proof. ∎
A.2 Morse index estimate
In the following we obtain a general estimate of the Morse index for the family of self-adjoint operators $(\mathcal{L}_{Z},D_{Z})$ in (5.1)-(1.7).
We start with the following proposition which provides a strategy for estimating the Morse index of self-adjoint extensions of a minimal symmetric operator (see [37], [44], Chapter X). We recall the notation $n_{\pm}(A)=\dim\ker(A^{*}\mp iI)$.

### v2

@@ 5.1 Morse index for $\mathcal{L}_{Z}$
        Suppose that $g_{{1}}\equiv 0$: then $f_{{1}}(-L)=f_{{1}}(L)=0$, $f^{\prime}_{{1}}(-L)=f^{\prime}_{{1}}(L)$, and so $f_{1}$ will be odd (see step (1) in item (A) above). Now, our single-lobe profile $\phi_{1}$ satisfies $\mathcal{L}_{1}\phi^{\prime}_{1}(x)=0$, for $x\in(-L,L)$, $\phi^{\prime}$ is odd, $\phi^{\prime}(x)>0$ for $x\in[-L,0)$, thus since $\lambda_{1}<0$ we obtain from the Sturm oscillation theorem that there is $r\in(-L,0)$ such that $\phi^{\prime}(r)=0$, it which is a contradiction. Then, $g_{1}$ is non-trivial.
@@
        Now suppose $g_{1}(0+)=0$: we consider the odd-extension $g_{1,\mathrm{odd}}\in H^{2}(\mathbb{R})$ of $g_{1}$ on all the line and the extension operator $\widetilde{\mathcal{L}}$ in (5.14) on the $\delta$-interaction domains $D_{\delta,\gamma}$ in (5.15). Then, $g_{1,\mathrm{odd}}\in D_{\delta,\gamma}$ for any $\gamma$ and by Perron-Frobenius property (Theorem A.5) we will obtain $n(\widetilde{\mathcal{L}})\geqq 2$, it which is a contradiction because $n(\widetilde{\mathcal{L}})\leqq 1$ for all $\gamma$ (see item (A)-(2) above).
@@
    Suppose $\lambda_{0}$ is a double eigenvalue: $\lambda_{0}=\lambda_{1}$. Then by Perron-Frobenius property in item (I)-(A) there holds $\xi_{\lambda_{0}},\zeta_{\lambda_{0}}>0$. Now, for $(f_{1},g_{1})\in D_{Z,0}$ an associated eigenfunction for $\lambda_{1}$, we have $(f_{1},g_{1})\perp(\xi_{\lambda_{0}},\zeta_{\lambda_{0}})$. Next, by item (I)-(B), Lemma 5.1 and $\lambda_{0}=\eta_{0}$, being $\eta_{0}$ the first eigenvalue (simple) for (PBP) in (5.6), we have that $f_{1}=r\xi_{\lambda_{0}}$ and $g_{1}=s\zeta_{\lambda_{0}}$ with $r=s\neq 0$ by continuity at zero. Hence, we arrive to a contradiction from the orthogonality property of the eigenfunctions. This finishes the proof of the Theorem.
@@
$\langle\mathcal{L}_{Z}(\phi_{1,k},\phi_{2,a})^{\top},(\phi_{1,k},\phi_{2,a})^{\top}\rangle=\int_{-L}^{L}\mathcal{T}(\phi_{1,k}(x))\phi_{1,k}(x)dx+\int_{L}^{\infty}\mathcal{T}(\phi_{2,a}(x))\phi_{2,a}(x)dx<0.$
@@
[…]
@@
In what follows we exactly determine which single-lobe kink state $(\phi_{1,k},\phi_{2,a})\in D_{Z}$, characterized by Propositions 4.7 and 4.11, will induce that the Morse index of $\mathcal{L}_{Z}$ is equal to one. As we will see the value of $L$, $k$, $c_{i}$ and the sign of $Z$ play a fundamental role.
@@
    with $a=a(k)$ determined by formula in (4.13). Then for $H(k)$ defined in (A.10) we have:
@@
    with $a=a(k)$ determined by formula in (4.13). Then, for $H(k)$ defined in (A.10) we have the following:
@@ 5.2 Trivial kernel for $\mathcal{L}_{Z}$
Let $Z\in(-\infty,\frac{2}{\pi c_{2}})$ be an admissible strength value to obtain a single-lobe kink state $(\phi_{1,k},\psi_{2,a})\in D_{Z}$, given by Propositions 4.7 and 4.11. Then, under the restrictions on $L,c_{1},c_{2},k$ in Theorem 5.10 and 5.12, we obtain $n(\mathcal{L}_{Z})=1$ and $\ker(\mathcal{L}_{Z})=\{\bf{0}\}$. In the case of a degenerate single-lobe kink state $(\pi,\phi_{2,0})\in D_{Z}$ with $Z=\frac{2}{\pi c_{2}}$, we also have that $\ker(\mathcal{L}_{Z})$ is trivial and $n(\mathcal{L}_{Z})=1$.
@@
    Suppose that $c=0$. This implies that $h\equiv 0$ and hence $f$ satisfies $\mathcal{L}_{1}f=0$ with Dirichlet-periodic boundary conditions on [L,L][-L,L], f(L)=f(L)=0f(L)=f(-L)=0 and $f^{\prime}(L)=f^{\prime}(-L)$. Suppose that $f\neq 0$:
@@
      Suppose that $f$ is even: then from oscillation results of the Floquet theory, we need to have that zero it is not the first eigenvalue for $\mathcal{L}_{1}$. Therefore, $f$ needs to change of sign and the number of zeros of $f$ on [L,L)[-L,L) is even. Hence $f$ has an odd number of zeros on (L,L)(-L,L) (as f(L)=0f(-L)=0). Then, $f(0)=0$ and $f^{\prime}(0)=0$. Hence, $f\equiv 0$.
@@
[…]
@@
    Next, we denote by $\{\eta_{n}\}$, $n\in\mathbb{N}_{0}$, the set of eigenvalues for the periodic-problem associated to $\mathcal{L}_{1}$ on [L,L][-L,L], and by $\{\mu_{n}\}$, $n\in\mathbb{N}_{0}$, the set of eigenvalues for the Dirichlet-problem associated to $\mathcal{L}_{1}$ on [L,L][-L,L]. Then, by classical Sturm-Liouville theory (see Theorem 4.8.1 in [41]) we have that $\eta_{0}$ is simple and, in particular, we have the following distribution of eigenvalues,
@@
      Suppose that $0>\mu_{1}$. We know that $\mathcal{L}_{1}\phi_{1}^{\prime}=0$, $\phi^{\prime}_{1}$ is odd and strictly decreasing on [L,0][-L,0]. Now, since the eigenfunction $\chi$ associated to $\mu_{1}$ is odd and $\chi(-L)=\chi(0)=0$, it follows from the Sturm Comparison Theorem (cf. [40, 41]) that $\phi^{\prime}_{1}$ needs to have one zero on (L,0)(-L,0). But this is impossible and we have a contradiction.
@@
      From the previous items (a) and (b) it follows that $\mu_{1}>0$ and, by (5.21) and Lemma 5.1 ($\mathcal{L}_{1}f=0$), we need to have $\eta_{1}=0$ and that the zero is simple. Moreover, the eigenfunction $f$ is even (for $f$ being odd, we have f(L)=f(L)=f(L)f(L)=-f(-L)=-f(L) and so $h(0)=f(L)=0$), which is a contradiction.
    Next, by Floquet theory, $f$ has exactly two different zeros $\beta,\beta$ ($\beta>0$) on (L,L)(-L,L). Therefore, $f(0)<0$, $f(\pm L)=h(0)>0$). Now, we consider the Wronskian function (constant) of $f$ and $\phi^{\prime}_{1}$ (we recall $\mathcal{L}_{1}f=\mathcal{L}_{1}\phi^{\prime}_{1}=0$),
@@
The statements (i)(iii)(i)-(iii) are immediate from the explicit formula for $\alpha_{a}$. This finishes the proof. ∎
@@
Under the restrictions on $Z,L,c_{1},c_{2}$ and $k$ given by Propositions 4.7 and 4.11 and by Theorems 5.10, 5.12 and 5.14, every member of the smooth family of single-lobe kink states, $k\to(\phi_{1,k},\phi_{2,a(k)})\in D_{Z}$, is spectrally and nonlinearly unstable for the sine-Gordon model (1.4) on a tadpole graph. The degenerate single-lobe kink state $(\pi,\phi_{2,0})\in D_{Z}$, with $Z=\frac{2}{\pi c_{2}}$ and
@@
Data availability
No data was used for the research described in the article.
@@ Acknowledgements
J. Angulo was supported in part by Universal/CNPq project and by CNPq/Brazil Grant. The work of R. G. Plaza was partially supported by DGAPA-UNAM, program PAPIIT, grant IN-104922 and by SECIHTI, grant CF-2023-G-122. R. G. Plaza is grateful to the IME-USP for their hospitality and to the FAPESP, Brazil (processo 2024/15816-0) for their financial support during a research stay in November 2024 when this work was carried out.
@@ Appendix A Appendix
In this section we formulate some tools from the extension theory of symmetric operators by Krein and von Neumann which are suitable for our needs (see, for instance, Naimark [29, 30] and Reed and Simon [36] for further information). In particular, we establish for $(\widetilde{\mathcal{L}},D_{\delta,\gamma})_{\gamma\in\mathbb{R}}$ in (5.14) and (5.15), a Perron-Frobenius result to be used in the accurate Morse index estimate $n(\mathcal{L}_{Z})\leqq 1$ for $(\mathcal{L}_{Z},D_{Z})$ in (5.1) and (1.7).
@@ A.1 Extension theory for the Laplacian operator on a tadpole graph
The following three results from the extension theory of symmetric operators are classical and can be found in [29, 30, 36].